Overhead Variances, Two- And Three-Variance Analyses

Oerstman, Inc., uses a standard costing system and develops its overhead rates from the current annual budget. The budget is based on an expected annual output of 124,000 units requiring 496,000 direct labor hours. (Practical capacity is 516,000 hours.) Annual budgeted overhead costs total $758,880, of which $555,520 is fixed overhead. A total of 119,100 units using 494,000 direct labor hours were produced during the year. Actual variable overhead costs for the year were $241,400, and actual fixed overhead costs were $555,450.

Required:

1. Compute overhead variances using a two-variance analysis.

Budget Variance $ Unfavorable
Volume Variance $ Unfavorable

2. Compute overhead variances using a three-variance analysis.

Spending Variance $ Unfavorable
Efficiency Variance $ Unfavorable
Volume Variance $
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Thanks!
Oerstman, Inc.
Overhead variances using a two-variance analysis Amount $ Note
Expected annual output 124,000.00 A
Direct labor hours required 496,000.00 B
Direct labor hours per unit                4.00 C=B/A
Annual budgeted overhead costs 758,880.00 D
Annual budgeted Fixed overhead 555,520.00 E
Annual budgeted Variable overhead 203,360.00 F=D-E
Variable overhead per hour                0.41 R=F/B
Variable Overhead Variance Finished Units Input per unit Input required Rate per unit Amount $
Standard Hours allowed for Actual Output at Standard Rate      119,100.00                  4.00       476,400.00                0.41 195,324.00 G
Actual Hours of Input, at Standard Rate       494,000.00                0.41 202,540.00 H
Actual Hours of Input, at Actual Rate 241,400.00 I
Variable overhead efficiency variance (H-G)        7,216.00 Unfavorable
Variable overhead Spending variance (I-H)     38,860.00 Unfavorable
Actual Fixed cost 555,450.00 P
Standard Fixed cost 555,520.00 J
Fixed Overhead Budget Variance             70.00 Unfavorable Q=J-P
Budget Variance
Variable overhead efficiency variance        7,216.00
Variable overhead Spending variance      38,860.00
Fixed Overhead Budget Variance             70.00
Overhead Budget Variance     46,146.00 Unfavorable
Fixed Overhead Volume Variance
Annual budgeted Fixed overhead 555,520.00 J
Annual budgeted (hours) 496,000.00 K
Standard Fixed overhead per hour                1.12 L=J/K
Standard Hours allowed for Actual Output 476,400.00 M
Applied fixed overhead 533,568.00 N=M*L
Overhead Volume Variance     21,952.00 Unfavorable O=J-N
Overhead variances using a three-variance analysis
Spending Variance
Variable overhead Spending variance      38,860.00
Fixed Overhead Budget Variance             70.00
Spending Variance     38,930.00 Unfavorable
Efficiency variance
Variable overhead efficiency variance        7,216.00 Unfavorable
Volume Variance
Fixed Overhead Volume Variance     21,952.00 Unfavorable